# First, Learn What You Don’t Know: Active Information Gathering for Driving at the Limits of Handling

### v1

@@ V Results
We validate the proposed active information gathering framework on a Toyota Supra. The states are estimated by an OxTS Inertial Navigation System [33], and the MPC problems are solved at $100\mathrm{Hz}$ on an on-board computer equipped with a 3.30GHz Intel Xeon E-2278GE CPU using a standard line search sequential quadratic programming method [34, Chapter 18] where quadratic programming subproblems are solved with the OSQP solver [35]. Reference trajectories for drifting are generated by the method in [36].
@@ V-A What do information-rich trajectories look like?
TABLE I: Covariance norms $\|\Lambda_{i}^{-1}\|_{2}$ for the prior model, the model adapted on the nominal OCP trajectory (Once-OCP), the model adapted on one Info-OCP trajectory (Once-Info-OCP), and the model adapted on two Info-OCP trajectories (Twice-Info-OCP).
$\|\Lambda_{i}^{-1}\|_{2}$ Prior Once-OCP Once-Info-OCP Twice-Info-OCP
@@
In Figure 4, we see that although the model is meta-trained only on drifting data, solutions to Info-OCP are not exactly drifting trajectories. Solutions to OCP roughly follow the straight-line reference trajectory, whereas the solution to Info-OCP applies larger engine torques and steers in each direction to excite the dynamics and provide informative data: Table I shows that following the solution to Info-OCP helps identify dynamics quicker than straight-line driving. Finally, after adapting on the first informative trajectory, the solution to Info-OCP applies even larger engine torques and steers the vehicle in each direction to reach larger yaw rates (Figure 4), which especially helps with reducing uncertainty over the yaw rate and sideslip dynamics (Table I). Intuitively, reaching higher yaw rates may help saturate tire forces and identify uncertain tire properties. Although solutions to Info-OCP are task-agnostic, identifying tire dynamics may be key to subsequent deployment in dynamic drifting tasks.
@@ V-B Information gathering enables reliable drifting
Next, we consider the task of drifting a Figure-8 trajectory, see Figure 1. This maneuver is challenging due to the fast and unstable transients during transitions between the circular arcs of the trajectory, shown in grey in Figure 6.
We evaluate the performance of three models for MPC: the posterior model adapted on the nominal trajectory from OCP (Once-OCP-Model); the posterior model adapted on the first information gathering trajectory from Info-OCP (Once-Info-Model); and the posterior model adapted on two information gathering trajectories (Twice-Info-Model). To assess the value of the proposed information gathering procedure in isolation, we do not perform online adaptation in this experiment. Each model is given three attempts to drift along the Figure-8, and we present the median tracking performance of each model in Figure 6.
@@
Performing at least one round of information gathering before deployment clearly helps in completing the first transition. We hypothesize that this is because adapting on trajectories from Info-OCP results in better estimates of yaw rate dynamics compared to adapting on the nominal OCP trajectory. Indeed, Figure 4 shows that the Info-OCP trajectories excite yaw rate dynamics more significantly than the OCP trajectory does. Moreover, Table I shows that adapting on Info-OCP trajectories results in smaller yaw rate covariance norms compared to adapting on the nominal OCP trajectory, suggesting more accurate yaw rate dynamics estimates after adapting on Info-OCP trajectories.
@@ V-C Online adaptation may not be enough for reliable control
Figure 7: Best performance in the tandem drifting maneuver of prior model with online adaptation and the model adapted on informative trajectories. (Top) Overhead view and (Bottom) time-series data of key signals. Online adaptation is disabled in the Twice-Info-Model.
Having established the utility of the information gathering and offline adaptation on trajectories generated via Info-OCP, we now evaluate the efficacy of online adaptation of the model through the last-layer update rule in (6) on a “tandem drifting” maneuver [^1] This maneuver was recently autonomously performed by two cars drifting in tandem   https://www.youtube.com/watch?v=8J_WiT-RD74 ., shown in Figure 7. This maneuver is challenging as the drift is initiated by the MPC controller, in contrast to the hand-brake initiation done in the Figure-8 maneuver. Thus, to successfully initiate the drift, it is essential that the model precisely describes the challenging initial transition between gripping and sliding.
We compare the performance of the prior model adapting online to the model previously adapted on the two information gathering trajectories from Section V-A. We test each of these models twice and report the best performance of each model in Figure 7.
Figure 7 shows that the model with online adaptation is able to refine its estimates online and enables accurate tracking of this challenging trajectory. However, while it successfully completes the maneuver and eventually yields accurate tracking, MPC with online adaptation alone leads to large initial errors, taking upward of $150$ meters to initiate the drift. This poor transient performance may be problematic for safety-critical applications of controllers for driving at the limits, e.g., for emergency obstacle avoidance [37].
In contrast, using the model that has adapted offline on the information gathering trajectories enables seamlessly drifting across all transitions, as shown in the position response and in the plots of the yaw rate and sideslip angle in Fig. 7. We postulate that the prior model does not capture the transition between gripping and sliding accurately enough to reliably initiate the drift. Although the model is meta-trained to be capable of rapid adaptation with little data, results show that online adaptation alone from the prior model does not suffice to seamlessly and quickly initiate the drift and leads to a large transient response. Information-rich data collection prior to deployment clearly helps in reliably performing this difficult maneuver while minimizing the transient response.
@@ VI Conclusion
We present an information gathering framework to actively and efficiently identify nonlinear dynamics for reliable subsequent deployment in MPC. Our approach hinges on an expressive vehicle dynamics model leveraging meta-learned features and Bayesian last-layer weight updates for rapid adaptation. Our results show that collecting informative data first unlocks reliable few-shot deployment within an MPC framework, and we expect that these findings translate to other challenging tasks that involve controlling unstable systems where small model errors may quickly result in large tracking errors. This methodology is especially valuable in applications where passive online adaptation may not be fast enough to correct the initial model mismatch in order to successfully complete the task, precisely because unstable systems are difficult to simultaneously learn and control.
In future work, the uncertainty estimates encoded in the Bayesian model could be used in the MPC formulation to robustify constraints during exploration and task deployment [20]. These uncertainty estimates could also be used to identify when a sufficient amount of data is collected or detect out-of-distribution behavior [38]. Finally, although the proposed method is task-agnostic and already demonstrates reliable performance in challenging drifting tasks, leveraging task-informed learning approaches [26, 14] may further enhance the performance of the method and further decrease the data requirements for active information gathering.

### v2

@@ V Results
We validate the proposed active information gathering framework on a Toyota Supra. The states are estimated at a rate of 100Hz by an OxTS Inertial Navigation System [35], and the MPC problems are solved at approximately $100\mathrm{Hz}$ on an on-board computer equipped with a 3.30GHz Intel Xeon E-2278GE CPU using a standard line search sequential quadratic programming method [36, Chapter 18] where quadratic programming subproblems are solved with the OSQP solver [37]. Reference trajectories for drifting are generated by the method in [38]. All videos of our experiments are available at https://tinyurl.com/d4pmh8kc.
@@ V-A What do information-rich trajectories look like?
TABLE I: Covariance norms $\|\Lambda_{i}^{-1}\|_{2}$ for the prior model, Once-OCP-Model (1-OCP-Model), Once-Info-Model (1-Info-Model), and Twice-Info-Model (2-Info-Model).
$\|\Lambda_{i}^{-1}\|_{2}$ Prior 1-OCP-Model 1-Info-Model 2-Info-Model
@@
In Fig. 4, we see that although the model is meta-trained only on drifting data, solutions to Info-OCP are not exactly drifting trajectories. Solutions to OCP roughly follow the straight-line reference trajectory, whereas the solution to Info-OCP applies larger engine torques and steers in each direction to excite the dynamics and provide informative data: Table I shows that following the solution to Info-OCP helps identify dynamics quicker than straight-line driving. Finally, after adapting on the first informative trajectory, the solution to Info-OCP applies even larger engine torques and steers the vehicle in each direction to reach larger yaw rates (see Fig. 4). This helps to reduce uncertainty over the yaw rate and sideslip dynamics (see Table I). Intuitively, reaching higher yaw rates may help saturate tire forces and identify uncertain tire properties. Although solutions to Info-OCP are task-agnostic, identifying tire dynamics may be key to subsequent deployment in dynamic drifting tasks.
@@ V-B Information gathering enables reliable drifting
Next, we consider the task of drifting a Figure-8 trajectory, see Fig. 1. This maneuver is challenging due to the fast and unstable transients during transitions between the circular arcs of the trajectory, shown in grey in Fig. 6.
We evaluate the performance of three models for MPC: Once-OCP-Model, Once-Info-Model, and Twice-Info-Model. To assess the value of the proposed information gathering procedure in isolation, we do not perform online adaptation in this experiment. Each model is given three attempts to drift along the Figure-8, and we present the median tracking performance of each model in Fig. 6 (additional statistics are reported in Appendix B). Tracking performance is measured in terms of the root mean squared lateral error along the reference.
@@
Performing at least one round of information gathering before deployment helps in completing the first transition. We hypothesize that this is because adapting on trajectories from Info-OCP results in better estimates of yaw rate dynamics compared to adapting on the nominal OCP trajectory. Indeed, Fig. 4 shows that the Info-OCP trajectories excite yaw rate dynamics more significantly than the OCP trajectory does. Moreover, Table I shows that adapting on Info-OCP trajectories results in smaller yaw rate covariance norms compared to adapting on the nominal OCP trajectory, suggesting more accurate yaw rate dynamics estimates after adapting on Info-OCP trajectories.
@@ V-C Online adaptation may not be enough for reliable control
Figure 7: Best performance in the loops maneuver for the prior model with online adaptation and the model adapted on informative trajectories. (Top) Overhead view and (Bottom) time-series data of key signals. Online adaptation is disabled in the Twice-Info-Model.
Having established the utility of the information gathering and offline adaptation on trajectories generated via Info-OCP, we now evaluate the efficacy of online adaptation of the model through the last-layer update rule in (6) on a loops maneuver shown in Fig. 7. This maneuver is challenging as the drift is initiated by the MPC controller, in contrast to the hand-brake initiation done in the Figure-8 maneuver. Thus, to successfully initiate the drift, it is essential that the model precisely describes the challenging initial transition between gripping and sliding.
We compare the performance of the prior model adapting online to the model previously adapted on the two information gathering trajectories from Sec. V-A. We test each of these models twice and report the best performance of each model in Fig. 7.
Fig. 7 shows that the model with online adaptation enables accurate tracking of this challenging trajectory. However, while it successfully completes the maneuver and eventually yields accurate tracking, MPC with online adaptation alone leads to large initial errors, requiring upward of $150$ meters to initiate the drift properly. This poor transient performance may be problematic for safety-critical applications of controllers for driving at the limits, e.g., for emergency obstacle avoidance [39].
In contrast, using the model that has adapted offline on the information gathering trajectories enables seamlessly drifting across all transitions, as shown in the in the plots of the tracking errors in position, yaw rate, and sideslip angle in Fig. 7. We postulate that the prior model does not capture the transition between gripping and sliding accurately enough to properly initiate the drift. Although the model is meta-trained to be capable of rapid adaptation with little data, results show that online adaptation alone from the prior model does not suffice to initiate the drift seamlessly and leads to a large transient response. Information-rich data collection prior to deployment clearly helps in reliably performing this difficult maneuver while minimizing the transient response.
@@ VI Conclusion
We present an information gathering framework to actively and efficiently identify nonlinear dynamics for reliable subsequent deployment in MPC. Our approach hinges on an expressive vehicle dynamics model leveraging meta-learned features and Bayesian last-layer weight updates for rapid adaptation. Our results show that collecting informative data first unlocks reliable few-shot deployment within an MPC framework, and we expect that these findings translate to other challenging tasks that involve controlling unstable systems where small model errors may result in large tracking errors. This methodology is especially valuable in applications where passive online adaptation may not be fast enough to correct the initial model mismatch to complete the task successfully, precisely because unstable systems are difficult to simultaneously learn and control.
In future work, the uncertainty estimates encoded in the Bayesian model could be used in the MPC formulation to robustify constraints during exploration and task deployment [20]. These uncertainty estimates could also be used to identify when a sufficient amount of data is collected or to detect out-of-distribution behavior [40]. Finally, although the proposed method is task-agnostic and already demonstrates reliable performance in challenging drifting tasks, leveraging control-oriented identification tools [28, 14, 41, 42] may further enhance the performance of the method and further decrease the data requirements for information gathering.
Appendix A Single-Track Bicycle Dynamics
The nominal dynamics model in (3) consists of a single-track bicycle model (see e.g. [5, 11]) given by
$\begin{bmatrix}\dot{r}\\[2.84526pt] \dot{v}\\ \\ \dot{\beta}\\ \\ \dot{\omega}_{r}\\[2.84526pt] \dot{e}\\ \Delta\dot{\varphi}\end{bmatrix}=\begin{bmatrix}(aF_{yf}\cos(\delta)+aF_{xf}\sin(\delta)-bF_{yr})/I_{z}\\ \big{(}{-}F_{yf}\sin(\delta-\beta)+F_{xf}\cos(\delta-\beta)+\\ \hskip 56.9055ptF_{yr}\sin(\beta)+F_{xr}\cos(\beta)\big{)}/m\\ -r+\big{(}F_{yf}\cos(\delta-\beta)+F_{xf}\sin(\delta-\beta)+\\ \hskip 56.9055ptF_{yr}\cos(\beta)-F_{xr}\sin(\beta)\big{)}/(mv)\\ (G\tau-F_{xr}R_{\textrm{w}})/I_{\textrm{w}}\\ v\sin(\Delta\varphi)\\ \dot{\beta}+r-\kappa_{\textrm{ref}}v\cos(\Delta\varphi)/(1-\kappa_{\textrm{ref}}e)\end{bmatrix},$
where $(a,b)$ are the distances from the center of gravity to the front and rear axles, respectively, $(I_{z},m)$ are the vehicle inertia and mass, respectively, $(R_{\text{w}},I_{\text{w}})$ are the wheel radius and the rear axle inertia, respectively, $G$ is the gear ratio from engine to rear axle, $\kappa_{\text{ref}}$ is the curvature of the reference trajectory, and $(F_{yf},F_{xf},F_{yr},F_{xr})$ denote tire forces and are modeled using an isotropic coupled slip brush Fiala model [43]. Following [5], since we test on a rear-wheel drive vehicle, we set $F_{xf}=0$.
Appendix B Additional Statistics on Figure-8 Trajectory
We report additional statistics to compare the performance of Once-OCP-Model, Once-Info-Model, and Twice-Info-Model for the task of drifting the Figure-8 trajectory from Sec. V-B. In Table II, we present the root mean square (RMS) lateral deviation $e$ over three trials for each of the three models. We also report the RMS sideslip angle error, $e_{\beta}=\beta-\beta_{\text{ref}},$ in Table III. Since neither Once-OCP-Model nor Once-Info-Model successfully completed the entirety of the Figure-8, we truncated the RMS computation before the spin out or the manual takeover for these models. Twice-Info-Model always achieves the lowest RMS errors and is also the only model that completes the entirety of the Figure-8 trajectory without ever spinning out.
TABLE II: Root mean squared (RMS) lateral deviation for Once-OCP-Model, Once-Info-Model, and Twice-Info-Model over three attempts of drifting the Figure-8 trajectory. Units are [m].
$\mathrm{RMS}(e)$ Once-OCP-Model Once-Info-Model Twice-Info-Model
Best 2.31 1.18 0.61
Median 3.19 1.31 0.83
Worst 7.35 3.64 1.29
TABLE III: Root mean squared (RMS) sideslip angle error $e_{\beta}=\beta-\beta_{\mathrm{ref}}$ for Once-OCP-Model, Once-Info-Model, and Twice-Info-Model over three attempts of drifting the Figure-8 trajectory. Units are [rad].
$\mathrm{RMS}(e_{\beta})$ Once-OCP-Model Once-Info-Model Twice-Info-Model
Best 0.137 0.105 0.068
Median 0.261 0.348 0.070
Worst 0.384 0.526 0.084